# RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives

### v3

@@ Abstract
Recent video generative models primarily rely on carefully written text prompts for specific tasks, like inpainting or style editing. They require labor-intensive textual descriptions for input videos, hindering their flexibility to adapt personal/raw videos to user specifications. This paper proposes RACCooN, a versatile and user-friendly video-to-paragraph-to-video generative framework that supports multiple video editing capabilities such as removal, addition, and modification, through a unified pipeline. RACCooN consists of two principal stages: Video-to-Paragraph (V2P) and Paragraph-to-Video (P2V). In the V2P stage, we automatically describe video scenes in well-structured natural language, capturing both the holistic context and focused object details. Subsequently, in the P2V stage, users can optionally refine these descriptions to guide the video diffusion model, enabling various modifications to the input video, such as removing, changing subjects, and/or adding new objects. The proposed approach stands out from other methods through several significant contributions: (1) RACCooN suggests a multi-granular spatiotemporal pooling strategy to generate well-structured video descriptions, capturing both the broad context and object details without requiring complex human annotations, simplifying precise video content editing based on text for users. (2) Our video generative model incorporates auto-generated narratives or instructions to enhance the quality and accuracy of the generated content. (3) RACCooN also plans to imagine new objects in a given video, so users simply prompt the model to receive a detailed video editing plan for complex video editing. The proposed framework demonstrates impressive versatile capabilities in video-to-paragraph generation (up to $9.4\%p\uparrow$ absolute improvement in human evaluations against the baseline), video content editing (relative $49.7\%\downarrow$ in FVD), and can be incorporated into other SoTA video generative models for further enhancement.
@@ 1 Introduction
Recent advances in video generative models (Yan et al., 2021, Hong et al., 2022, Esser et al., 2023, Mei & Patel, 2023, Chai et al., 2023, Ceylan et al., 2023, Blattmann et al., 2023, Wu et al., 2023a), including Sora (openai, 2024), have demonstrated remarkable capabilities in creating high-quality videos. Simultaneously, video editing models (Geyer et al., 2023, Qi et al., 2023, Wang et al., 2024b, Yu et al., 2023, Wu et al., 2024, Zhang et al., 2023b) have gained significant attention, thanks to their promising applications that allow users to modify videos according to user-written textual instructions, effectively altering video content and attributes. Despite these advancements, significant challenges remain in developing a versatile and user-friendly framework that facilitates easy video modification for personal use. The primary challenges include: 1) the complexity of training a unified framework encompassing multiple video editing skills (e.g., remove, add, or change an object). Training a single model to perform various editing skills is highly challenging, and recent video editing methods often focus on specific tasks, such as background inpainting (Yu et al., 2023, Wu et al., 2024), or attribute editing (Geyer et al., 2023, Qi et al., 2023, Jeong & Ye, 2023). 2) the necessity for well-structured textual prompts that accurately describe videos and can be edited to support diverse video editing skills. The quality of prompts critically influences the models’ capabilities and the quality of their outputs. Generating detailed prompts is time-consuming and costly, and the quality varies depending on the expertise of the annotators. Although Multimodal Large Language Models (MLLMs) (Liu et al., 2023b, Munasinghe et al., 2023, Yang et al., 2023, Yu et al., 2024) have been explored for automatically describing videos, they often overlook critical details in complex scenes. This oversight compromises the development of a seamless pipeline, hindering both user convenience and the effectiveness of video generative models.
@@
To tackle these limitations, as shown in Fig. 1, we introduce RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives, a novel video-to-paragraph-to-video (V2P2V) generative framework that facilitates diverse video editing (Remove, Add, and Change) capabilities based on auto-generated narratives. RACCooN allows for the seamless removal and modification of subject attributes, as well as the addition of new objects to videos without requiring densely annotated video prompts or extensive user planning. Our framework operates in two main stages: video-to-paragraph (V2P) and paragraph-to-video (P2V). In the V2P stage, we introduce a new video descriptive framework built on a pre-trained Video-LLM backbone (PG-Video-LLaVA (Munasinghe et al., 2023)). We find that existing Video-LLMs effectively capture holistic video features, yet often overlook detailed cues that are critical for accurate video editing, as users may be interested in altering these missing contexts. To address this, we propose a novel multi-granular video perception strategy that leverages superpixels (Li & Chen, 2015, Ke et al., 2023) to capture diverse and informative localized contexts throughout a video. We first extract fine-grained superpixels using a lightweight predictor (Yang et al., 2020) and then apply overlapping k-means clustering (Cleuziou, 2007, Whang et al., 2015, Khanmohammadi et al., 2017) to segment visual scenes into various levels of granularity. The suggested localized spatiotemporal segmentation assists the LLM’s comprehension of objects, actions, and events within the video, enabling it to generate fluent and detailed natural language descriptions. Next, in the P2V stage, to integrate multiple editing capabilities into a single model, we fine-tuned a video inpainting model that can paint video objects accurately with detailed text, object masks, and condition video. Then, by utilizing user-modified prompts from generated descriptions in the V2P stage, our video diffusion model can accurately paint corresponding video regions, ensuring that textual updates from prompts are reflected in various editing tasks. Moreover, to better support our model training, we have collected the Video Paragraph with Localized Mask (VPLM) dataset—a collection of over 7.2K high-quality video-paragraph descriptions and 5.5k detailed object descriptions with masks, annotated from the publicly available dataset using GPT-4V (Achiam et al., 2023).
@@
We emphasize that RACCooN enhances the quality and versatility of video editing by leveraging detailed, automatically generated textual prompts that minimize ambiguity and refine the scope of generation. We validate the extensive capabilities of the RACCooN framework in both V2P generation, text-based video content editing, and video generation on ActivityNet (Krishna et al., 2017), YouCook2 (Zhou et al., 2018a), UCF101 (Soomro et al., 2012), DAVIS (Pont-Tuset et al., 2017), and our proposed VPLM datasets. On the V2P side, RACCooN outperforms several strong video captioning baselines (Li et al., 2023, Munasinghe et al., 2023, Liu et al., 2023b), particularly improving by average +9.1%p on VPLM and up to +9.4%p on YouCook2 compared to PG-VL (Munasinghe et al., 2023), based on both automatic metrics and human evaluation. On the P2V side, RACCooN surpasses previous strong video editing/inpainting baselines (Geyer et al., 2023, Qi et al., 2023, Wang et al., 2024b, Yu et al., 2023, Wu et al., 2024) over three subtasks of video content editing (remove, add, and change video objects) over 9 metrics. We also demonstrate that the proposed RACCooN framework can enhance SoTA video generative models by leveraging detailed auto-generated textual prompts. We further conduct extensive ablation and visualizations to validate the improvement quantitatively and qualitatively. Our contributions are as follows:
@@
    Framework Contribution: RACCooN offers a user-friendly and unified framework for various video editing tasks. It provides improved interpretability and interactive experiences by automatically generating detailed, object-centric video descriptions and layout plans tailored to different editing objectives, which cannot be done through a simple combination of existing models.
@@
    Technical Contribution: We present a novel multi-granular pooling strategy to capture local video contexts, enhancing video comprehension by generating fluent and detailed descriptions in a zero-shot setting. This enables users to create new videos that retain the visual characteristics of the input and focus on specific context editing.
@@
    Training/Dataset Contribution: To enable RACCooN to follow complex and varied user requests for video editing, we present the VPLM dataset, which contains 7.2K high-quality detailed video paragraphs and 5.5K object-level detailed caption-mask pairs. This dataset facilitates accurate V2P and P2V stages with high-quality, localized textural prompts and videos.
@@ 2 Related Work
The recent trend in video-language tasks focuses on generating comprehensive textual descriptions for long and complex video content (Shen et al., 2017, Krishna et al., 2017, Wang et al., 2018, Tewel et al., 2022, Wu et al., 2023b). Vid2Seq (Yang et al., 2023) introduces a novel dense event captioning approach for narrated videos, with time tokens and event boundaries. Video-LLaVA variants (Lin et al., 2023a, Munasinghe et al., 2023) present a large multimodal model integrating text, video, and audio inputs for generative and question-answering tasks. Similarly, LLaVA-Next (Zhang et al., 2024) improves zero-shot video understanding by transferring multi-image knowledge through concatenated visual tokens. While these methods are effective in video description, they often miss key contextual details (Zhang et al., 2023a, Li et al., 2023). Our RACCooN captures both holistic and localized details by leveraging localized spatiotemporal information, enhancing video editing and generation capabilities.
Prompt-to-Video Editing.
@@
Video editing (Ceylan et al., 2023, Liu et al., 2023c, Couairon et al., 2023, Kondratyuk et al., 2023, Wang et al., 2023, Zhang et al., 2023c) involves enhancing, modifying, or manipulating video content for desired effects. VideoComposer (Wang et al., 2024b) offers a multi-source controllable video generative framework. TokenFlow (Geyer et al., 2023) adapts text-to-image diffusion with flow matching for consistent text-driven video editing. LGVI (Wu et al., 2024) integrates an MLLM for complex language-based video inpainting. These methods often focus on specific tasks and may inadvertently alter unrelated regions due to limited contextual information. Our V2P2V framework overcomes these limitations by using auto-generated, detailed descriptions to integrate key contexts into diverse editing tasks.
@@
3 RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives
Figure 2: Illustration of RACCooN framework. RACCooN generates video descriptions with the three distinct pooled visual tokens, including Multi-Granular Spatiotemporal (MGS) Pooling. Next, users can edit the generated descriptions by adding, removing, or modifying words to create new videos. Note that for adding object tasks, if users do not provide layout information for the objects they want to add, RACCooN can predict the target layout in each frame.
Conditional video generation and editing models, designed to accurately reflect user requirements described in the text, often face challenges with complex scenes due to limited, high-level descriptions and a lack of detailed video understanding. Despite improvements from recent advances in MLLMs, these models still struggle to capture complex spatial-temporal dynamics, often omitting crucial objects and details. Training a text-to-video model with such vague prompts compromises output specificity, and leads the model to generate average, arbitrary content that fails to capture user instructions’ nuances. To address these issues, we introduce RACCooN, a user-friendly, two-stage video-to-paragraph-to-video editing framework. Initially, RACCooN generates detailed, structured paragraphs from videos, capturing holistic content and key local objects through multi-granular spatiotemporal information. These detailed descriptions are then used for conditional video generation and editing, enabling users to add, remove, or change video objects directly by interacting with the generated descriptions, thus enhancing the specificity and relevance of the output.
@@ 3 RACCooN: A Versatile Instructional Video Editing Framework with Auto-Generated Narratives
3.1 V2P: Auto-Descriptive Framework with Multi-Granular Video Perception
@@
Multimodal LLM for Video Paragraph Generation. In the V2P stage, the RACCooN framework generates well-structured, detailed descriptions for both holistic videos and local objects. It employs a multimodal LLM with three main components: a visual encoder $E$, a multimodal projector, and an LLM. Given an input video $\bm{x}\in\mathbb{R}^{F\times C\times H\times W}$, where $F$, $C$, $H$, and $W$ represent the number of frames, channels, height, and width, respectively, we extract video features using the visual encoder: $\bm{e}=E(\bm{x})\in\mathbb{R}^{t\times h\times w\times d}$. Here, $t$, $h$, $w$, and $d$ denote the encoded temporal dimension, the height and width of the tokens, and the feature dimension. To understand complex videos with multiple scenes, we use three pooling strategies: spatial pooling, temporal pooling, and multi-granular spatiotemporal pooling. Spatial pooling $\bm{e}^{s}=\text{Pooling}^{s}(\bm{e})\in\mathbb{R}^{t\times d}$ aggregates tokens within the same frame, while temporal pooling $\bm{e}^{t}=\text{Pooling}^{t}(\bm{e})\in\mathbb{R}^{(h\cdot w)\times d}$ averages features across the temporal dimension for the same region. Despite these strategies helping the LLM grasp the video holistically in space or time, they often overlook capturing key objects or actions localized throughout the video stream, especially in untrimmed, dynamic, multi-scene videos.
@@
Multi-Granular Spatiotemporal Pooling. To address this issue, we introduce a novel superpixel-based spatiotemporal pooling strategy, coined multi-granular spatiotemporal pooling (MGS pooling). As illustrated in Fig. 2 left top, this strategy is designed to capture localized information via superpixels across spatial and temporal dimensions. Superpixels (Li & Chen, 2015, Giordano et al., 2015, Yang et al., 2020, Ke et al., 2023) are small and coherent clusters of pixels that share similar characteristics, such as color or texture. These clusters provide an efficient representation of visual scenes and are resilient to frame noise since they average out the pixel values within each cluster, effectively smoothing out variations induced by noise. As shown in Fig. 3, we use a lightweight superpixel predictor $\sigma(\cdot)$ (Yang et al., 2020) to generate superpixels across video frames, capturing the granular visuality of each local area. However, due to their limited coverage area, these fine-grained visual features often fail to capture attribute-level semantics, such as objects and actions (Zhang et al., 2023a, Li et al., 2023). Motivated by the importance of varying the compositions of multiple superpixels for different contexts in video understanding, we propose the use of overlapping k-means clustering (Cleuziou, 2007, Whang et al., 2015) for the obtained video superpixels, which improves the granularity from fine to coarse. This approach allows the LLM to gather informative cues about various objects and actions. We first obtain the pixel features and the superpixel index vector for the video pixels: $\bm{S},\bm{g}=\sigma(\bm{x},\bm{g}_{\text{init}})$, where $\bm{g}_{\text{init}}$ is the input superpixel indices, initialized by a region-based grid. Given the averaged pixel features of each superpixel, $\bar{\bm{S}}\in\mathbb{R}^{|\bm{g}|\times d_{p}}$, where $d_{s}$ denotes the pixel feature size, we generate the MGS tokens $\bm{e}^{l}$:
@@
$\begin{split}\widehat{\bm{e}}&=\textsc{concat}[\bm{e}^{s};\bm{e}^{l};\bm{e}^{t}]\cdot\bm{W}^{\top},\\ \bm{a}&=LLM(\textsc{concat}[\bm{e}^{p};\widehat{\bm{e}}]),\end{split}$ (2)
@@
where $\bm{W}\in\mathbb{R}^{d\times d^{\prime}}$ is the weight matrix for linear projection into the text embedding dimension $d^{\prime}$. We highlight that our video description framework serves as an integrated, user-interactive tool for video-to-paragraph generation and video content editing. (Fig. 2 top right).
@@
3.2 P2V: User-Interactive Video Content Editing with Automatically Generated Descriptions

### v4

# RACCooN: Versatile Instructional Video Editing with Auto-Generated Narratives

@@ Abstract
Recent video generative models primarily rely on detailed, labor-intensive text prompts for tasks, like inpainting or style editing, limiting adaptability for personal/raw videos. This paper proposes RACCooN, a versatile and user-friendly video-to-paragraph-to-video editing method, supporting diverse video editing capabilities, such as removal, addition, and modification, through a unified pipeline. RACCooN consists of two main stages: Video-to-Paragraph (V2P), which automatically generates structured descriptions of scene and object details, and Paragraph-to-Video (P2V), where users can refine these to guide a video diffusion model for flexible content edits, including removing, changing, or adding objects. Key contributions of RACCooN include: (1) A multi-granular spatiotemporal pooling strategy for structured video understanding, capturing both global context and fine-grained object details to enable precise text-based video editing without complex human annotations. (2) A video generative model fine-tuned on a curated video-paragraph-mask dataset for improved editing and inpainting. (3) The ability to generate new objects by forecasting motion via auto-generated mask planning. In the end, users can easily edit complex videos with RACCooN’s automatic explanations and guidance. We demonstrate its versatile capabilities in video-to-paragraph generation (up to $9.4\%p\uparrow$ improvement in human evaluations), video content editing (relative $49.7\%\downarrow$ in FVD).
RACCooN: Versatile Instructional Video Editing
with Auto-Generated Narratives
Jaehong Yoon1,2,thanks: Equal Contribution.    Shoubin Yu1,11footnotemark: 1     Mohit Bansal1 1UNC Chapel Hill   2Nanyang Technological University jaehong.yoon@ntu.edu.sg  {shoubin, mbansal}@cs.unc.edu https://raccoon-mllm-gen.github.io/
@@ 1 Introduction
Recent advances in video generative models (Yan et al., 2021; Hong et al., 2022; Esser et al., 2023; Mei and Patel, 2023; Chai et al., 2023; Blattmann et al., 2023; Wu et al., 2023a), including Sora (openai, 2024), have demonstrated remarkable capabilities in creating high-quality videos. Simultaneously, video editing models (Geyer et al., 2023; Qi et al., 2023; Wang et al., 2024b; Yu et al., 2023; Wu et al., 2024; Zhang et al., 2023b) have gained significant attention for enabling users to modify content using user-written instructions. However, building a versatile, user-friendly framework that facilitates easy video modification for personal use remains challenging. Key challenges are as follows:
1) Training a unified model for multiple editing tasks (e.g., add, remove, change objects) is difficult Yu et al. (2025), and most methods focus on narrow tasks, such as background inpainting (Yu et al., 2023; Wu et al., 2024), or attribute editing (Geyer et al., 2023; Qi et al., 2023; Jeong and Ye, 2023).
2) The need for well-structured prompts that accurately describe videos and support diverse editing tasks is critical, as prompt quality directly impacts model performance. However, generating such prompts is costly and time-consuming, with quality varying by annotator expertise. While Multimodal Large Language Models (MLLMs) (Liu et al., 2023b; Yang et al., 2023; Yu et al., 2024) have been explored for automatic video description, they often miss key details in complex scenes, limiting seamless and user-friendly editing.
@@
To tackle these limitations, we introduce RACCooN: A Versatile Instructional Video Editing with Auto-Generated Narratives, a novel video-to-paragraph-to-video (V2P2V) generative method that facilitates diverse video editing (Remove, Add, and Change) capabilities based on auto-generated video descriptions. RACCooN enables seamless removal or modification of subject attributes and addition of new objects in videos without the need for densely annotated prompts or extensive user input. RACCooN operates in two main stages: video-to-paragraph (V2P) and paragraph-to-video (P2V). In the V2P stage, we introduce a new video descriptive module built on a pre-trained Video-LLM backbone (PG-Video-LLaVA (Munasinghe et al., 2023)). We find that existing Video-LLMs effectively capture holistic video features, yet often overlook detailed cues that are critical for accurate video editing, as users may be interested in altering these missing contexts. To address this, we propose a novel multi-granular video perception strategy that leverages superpixels (Li and Chen, 2015; Ke et al., 2023) to capture diverse and informative localized contexts throughout a video. We first extract fine-grained superpixels using a lightweight predictor (Yang et al., 2020) and then apply overlapping k-means clustering (Cleuziou, 2007; Whang et al., 2015) to segment visual scenes into various levels of granularity. The suggested localized spatiotemporal segmentation assists the LLM’s comprehension of objects, actions, and events within the video, enabling it to generate fluent and detailed natural language descriptions. Next, in the P2V stage, we fine-tune a video inpainting model to support multiple editing tasks using auto-generated detailed descriptions and object masks. Then, by leveraging user prompts derived from generated descriptions in the V2P stage, our video diffusion model accurately paint corresponding video regions, ensuring that textual edits are faithfully reflected across various editing tasks.
To further support model training, we introduce the Video Paragraph with Localized Mask (VPLM) dataset comprising over 7.2K quality video-paragraph data and 5.5K detailed object-level captions with masks, annotated from public datasets using GPT-4V (Achiam et al., 2023).
@@
We emphasize that RACCooN enhances the quality and versatility of video editing by leveraging detailed, automatically generated prompts that minimize ambiguity and refine the scope of generation. We validate the extensive capabilities of the RACCooN in both V2P generation, text-based video content editing, and video generation on ActivityNet (Krishna et al., 2017), YouCook2 (Zhou et al., 2018a), UCF101 (Soomro et al., 2012), DAVIS (Pont-Tuset et al., 2017), and our proposed VPLM datasets. On the V2P side, RACCooN outperforms several strong video captioning baselines (Li et al., 2023; Munasinghe et al., 2023; Liu et al., 2023b), particularly improving by average +9.1%p on VPLM and up to +9.4%p on YouCook2 compared to PG-VL (Munasinghe et al., 2023), based on both automatic metrics and human evaluation. On the P2V side, RACCooN surpasses previous strong video editing/inpainting baselines (Geyer et al., 2023; Qi et al., 2023; Wang et al., 2024b; Yu et al., 2023; Wu et al., 2024) over three subtasks of video content editing (remove, add, and change video objects) over 9 metrics. Our contributions are as follows:
@@
    Framework Contribution: RACCooN offers a user-friendly, unified framework for diverse video editing tasks, enhancing interpretability and interaction by generating detailed, object-centric descriptions and layout plans tailored to editing goals, surpassing existing model combinations.
@@
    Technical Contribution: We present a novel multi-granular pooling strategy to capture local video contexts, enhancing video comprehension by generating fluent and detailed descriptions in a zero-shot setting. This enables users to create new videos that retain the visual characteristics of the input and support targeted context editing.
@@
    Training/Dataset Contribution: To enable RACCooN to handle complex and diverse video editing requests, we present the VPLM dataset, comprising 7.2K high-quality video paragraphs and 5.5K object-level caption-mask pairs. These well-structured annotations enable accurate V2P and P2V stages at both video and object levels.
Figure 2: Illustration of RACCooN. RACCooN generates video descriptions with the three distinct pooled visual tokens, including Multi-Granular Spatiotemporal (MGS) Pooling. Next, users can edit the generated descriptions by adding, removing, or modifying words to create new videos. Note that for adding object tasks, if users do not provide layout information for the objects they want to add, RACCooN can predict the target layout in each frame.
@@ 2 Related Work
Video-to-Paragraph Generation. Recent video-language tasks focus on generating comprehensive textual descriptions for long and complex video content (Shen et al., 2017; Krishna et al., 2017; Wang et al., 2018; Tewel et al., 2022; Wu et al., 2023b). Vid2Seq (Yang et al., 2023) introduces a novel dense event captioning approach for narrated videos, with time tokens and event boundaries. Video-LLaVA variants (Lin et al., 2023a; Munasinghe et al., 2023) present a large multimodal model integrating text, video, and audio inputs for generative and question-answering tasks. Similarly, LLaVA-Next (Zhang et al., 2024) improves zero-shot video understanding by transferring multi-image knowledge through concatenated visual tokens. While these methods are effective in video description, they often miss key contextual details (Zhang et al., 2023a; Li et al., 2023). RACCooN captures both holistic and object-level details by leveraging localized spatiotemporal information, enhancing video editing and generation.
@@
Prompt-to-Video Editing. Video editing (Ceylan et al., 2023; Liu et al., 2023c; Couairon et al., 2023; Kondratyuk et al., 2023; Wang et al., 2023; Zhang et al., 2023c) involves enhancing, modifying, or manipulating video content for desired effects. VideoComposer (Wang et al., 2024b) offers a multi-source controllable video generative framework. TokenFlow (Geyer et al., 2023) adapts text-to-image diffusion with flow matching for consistent text-driven video editing. LGVI (Wu et al., 2024) integrates an MLLM for complex language-based video inpainting. These methods often focus on specific tasks and may inadvertently alter unrelated regions due to limited contextual information. Our V2P2V framework overcomes these limitations by using auto-generated, detailed descriptions to integrate key contexts into diverse editing tasks.
@@
3 A Versatile Instructional Video Editing with Auto-Generated Narratives
Conditional video generation and editing models struggle with complex scenes due to vague text descriptions and limited video understanding. To address this, we introduce RACCooN, a user-friendly, two-stage video-to-paragraph-to-video editing approach, with each stage detailed in Sec.˜3.1 and Sec.˜3.2. We also introduce the VPLM dataset, specifically curated to train models for detailed video editing, along with the training pipeline of RACCooN, detailed in Sec.˜3.3.
@@ 3 A Versatile Instructional Video Editing with Auto-Generated Narratives
3.1 V2P: Auto-Descriptive Module with Multi-Granular Video Perception
@@
Multimodal LLM for Video Paragraph Generation. In the V2P stage, the RACCooN generates well-structured, detailed descriptions for both holistic videos and local objects. It employs a multimodal LLM with three main components: a visual encoder $E$, a multimodal projector, and an LLM. Given an input video $\bm{x}\in\mathbb{R}^{F\times C\times H\times W}$, where $F$, $C$, $H$, and $W$ represent the number of frames, channels, height, and width, respectively, we extract video features using the visual encoder: $\bm{e}=E(\bm{x})\in\mathbb{R}^{t\times h\times w\times d}$. Here, $t$, $h$, $w$, and $d$ denote the encoded temporal dimension, the height and width of the tokens, and the feature dimension. To understand complex videos with multiple scenes, we use three pooling strategies: spatial pooling, temporal pooling, and multi-granular spatiotemporal pooling. Spatial pooling $\bm{e}^{s}=\text{Pooling}^{s}(\bm{e})\in\mathbb{R}^{t\times d}$ aggregates tokens within the same frame, while temporal pooling $\bm{e}^{t}=\text{Pooling}^{t}(\bm{e})\in\mathbb{R}^{(h\cdot w)\times d}$ averages features across the temporal dimension for the same region. Despite these strategies helping the LLM grasp the video holistically in space or time, they often overlook capturing key objects or actions localized throughout the video stream, especially in untrimmed, dynamic, multi-scene videos.
@@
Multi-Granular Spatiotemporal Pooling. To address this issue, we introduce a novel superpixel-based spatiotemporal pooling strategy, coined multi-granular spatiotemporal pooling (MGS pooling). As illustrated in Fig.˜2 left top, this strategy is designed to capture localized information via superpixels across spatial and temporal dimensions. Superpixels (Li and Chen, 2015; Yang et al., 2020; Ke et al., 2023) are small and coherent clusters of pixels that share similar characteristics, such as color or texture. These clusters provide an efficient representation of visual scenes and are resilient to frame noise since they average out the pixel values within each cluster, effectively smoothing out variations induced by noise. As shown in Fig.˜3, we use a lightweight superpixel predictor $\sigma(\cdot)$ (Yang et al., 2020) to generate superpixels across video frames, capturing the granular visuality of each local area. However, due to their limited coverage area, these fine-grained visual features often fail to capture attribute-level semantics, such as objects and actions (Zhang et al., 2023a; Li et al., 2023). Motivated by the importance of varying the compositions of multiple superpixels for different contexts in video understanding, we propose the use of overlapping k-means clustering (Cleuziou, 2007; Whang et al., 2015) for the obtained video superpixels, which improves the granularity from fine to coarse. This approach allows the LLM to gather informative cues about various objects and actions. We first obtain the pixel features and the superpixel index vector for the video pixels: $\bm{S},\bm{g}=\sigma(\bm{x},\bm{g}_{\text{init}})$, where $\bm{g}_{\text{init}}$ is the input superpixel indices, initialized by a region-based grid. Given the averaged pixel features of each superpixel, $\bar{\bm{S}}\in\mathbb{R}^{|\bm{g}|\times d_{p}}$, where $d_{p}$ denotes the pixel feature size, we generate the MGS tokens $\bm{e}^{l}$:
@@
[…]
@@
where $\bm{W}\in\mathbb{R}^{d\times d^{\prime}}$ is the weight matrix for linear projection into the text embedding dimension $d^{\prime}$. We highlight that our video description module serves as an integrated, user-interactive tool for video-to-paragraph generation and video content editing. (Fig.˜2 top right).
@@
3.2 P2V: User-Interactive Video Editing with Auto-Generated Descriptions